A survey of magazine subscribers showed that 45.9% rented a car during the past 12 months for business reasons, 52% rented a car during the past 12 months for personal reasons, and 40% rented a car during the past 12 months for both business and personal reasons.

Round your answers to three decimal places.

a. What is the probability that a subscriber rented a car during the past 12 months for business or personal reasons?

b. What is the probability that a subscriber did not rent a car during the past 12 months for either business or personal reasons?
Answer #1

a) P(Rented for Business or Personal reasons) = P(Business) + P(Personal) - P(Both) = 0.459 + 0.52 - 0.40 = 0.471

b) P(Not rented for either of the reasons) = 1 - 0.471 = 0.529